HUBBLE SPACE TELESCOPE DAILY REPORT #2472 PERIOD COVERED: 0000Z (UTC) 10/05/99 - 0000Z (UTC) 10/06/99 Daily Status Report as of 279/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 8409 (STIS CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and to track the evolution of hot columns. The observations were executed as scheduled, and no anomalies were reported. 1.2 Completed Two Sets of WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no problems were reported. 1.3 Completed Two Sets of STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.4 Completed WF/PC-2 8216 (T Tauri Star Snapshot Survey II. Completing the CTTs Sample) The WF/PC-2 was used to make observations of the T-Tauri class star known as V955-Tau as part of Snapshot Survey II. This database will be of fundamental importance to understanding the nature of planet-forming environments and their frequency in the galaxy. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of WF/PC-2 8457 (UV Earthflats) The WF/PC-2 was used to make a ultraviolet Earth flat calibration image. This is to improve the quality of pipeline the ultraviolet flatfields. The observations were completed as scheduled, and no problems were reported. 1.6 Completed WF/PC-2 8192 (A Snapshot Survey of Probable Nearby Galaxies) The WF/PC-2 was used to make a snapshot survey images of a probable nearby galaxy Cam-A. The census of galaxies in the Local Volume of space is very incomplete, particularly at the faint end. The observations were executed as scheduled, and no anomalies were noted 1.7 Completed FGS 8309 (A Link Between Massive Binary Stars and Non-thermal Radio Emission) The Fine Guidance Sensors were used to make astrometric observations of the massive Wolf-Rayet binary star WR40 and several nearby reference stars. Non-thermal radio emission in Wolf-Rayet stars is explained in terms of synchrotron emission from shocks in the stellar wind. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Two Sets of STIS/MA1 8426 (Cycle 8 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a routine Cycle-8 MAMA dark noise calibration measurement. This proposal will provide the primary means of checking on the health of the MAMA detectors. This is done through frequent monitoring of the background count rate. The observations were completed as planned, and no problems were reported. 1.9 Completed Two Sets of STIS/CCD 8406 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition V proposal. The observations were completed as scheduled, and no anomalies were reported. 1.10 Completed STIS/MA2 8275 (Ultraviolet Snapshots of 3CR Radio Source Counterparts at Low Z) The Space Telescope Imaging Spectrograph (MA2) was used to make ultraviolet snapshot observations of the radio galaxy 3C236.0. Radio galaxies are an important class of extragalactic objects, representing one of the most energetic astrophysical phenomena, while extraordinarily useful as probes of their environments and of the early Universe. The observations were completed as scheduled, and no anomalies were reported. 1.11 Completed STIS/CCD/MA1/MA2 8224 (UV Imaging of Europa & Ganymede: Unveiling Satellite Aurora & Electrodynamical Interactions) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make ultraviolet observations of the aurora of Jupiter’s moon Europa. This is to isolate atomic oxygen {1304 and 1356 Angstrom} and hydrogen {Lyman-Alpha} emissions, to study the interaction of the Jovian magnetosphere with its moons. A wavelength calibration image was also taken. As described in HSTAR 7295 and yesterday’s report, the acquisition for this proposal defaulted to fine lock backup on one FGS only, possibly affecting 15 observations. Later in the proposal, the losses of lock mentioned in 2.1 and HSTAR 7296 occurred, affecting one observation. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.12 Completed WF/PC-2 8441 (WF/PC-2 Decontaminations and Associated Observations Pt. 1/2) The WF/PC-2 was used to make its monthly decontamination operations. Also included are instrument monitors tied to decons: photometric stability check, focus monitor, pre- and post-decon internals {bias, internal flats, k-spots, and darks}, and ultraviolet throughput checks. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed STIS/CCD 8236 (Black Holes and Gas Disks in a Complete Sample of Radio-Loud Ellipticals - II: Kinematics) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the radio-loud elliptical galaxy NGC-541. From these observations we will measure the black hole masses, determine the nature, kinematics, and structure of the central gas disks and obtain a modeled upper limit to their mass accretion rates. A wavelength calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.14 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed WF/PC-2 8181 (Resolving Sirius-like Binaries) The WF/PC-2 was used to make observations of the star HD156638 in order to try and resolve these Sirius-like binary star systems. Subsequent determinations of the visual orbits of the resolved binaries with the HST will provide dynamical measurements of the white-dwarf masses. The observations were completed as scheduled, and no anomalies were reported. 1.16 Completed WF/PC-2 7374 (Probing Low Density Cluster Environments at Moderate-to-High Redshifts) The WF/PC-2 was used to make observations of the galactic cluster Gal-Clus-0273. Recent HST WF/PC-2 images of rich clusters at moderate-to-high redshifts have revealed that morphological information is a crucial ingredient in understanding the physical processes responsible for the evolution of the galaxy population in clusters. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed STIS/CCD 8549 (POMS Test Proposal: STIS non-scripted parallel proposal continuation III) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations as part of a non-scripted POMS Test Proposal. The observations were completed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 7 Successful: 7 Per HSTAR 7296, there were losses of lock at 278/102255Z and at 278/104354Z (this followed the acquisition default to control on one FGS described in yesterday’s report and HSTAR 7295). The take data flag remained down until shortly after the last loss. The proposal described in 1.11 was affected. Scheduled Re-acquisitions: 8 Successful: 8 2.2 FHST Updates: Scheduled: 17 Successful: 17 2.3 Operations Notes: The SSR EDAC error counter was cleared twice per ROP SR-1. The MA transmitter 1 and 2 RF and REFL power parameters went out-of-limits at 278/112819Z (for 26 minutes) and at 278/120607Z (for 10 minutes) respectively. HSTAR 7297 was written. At approximately 278/2008Z, several EPS voltage mnemonics flagged out-of-limits low for approximately 20 seconds each as the HST approached the end of orbit night. HSTAR 7299 was written. As directed by COP 18.24, the STIS memory was dumped at 279/0450Z. This followed an unsuccessful attempt at 278/110759Z when a missing “tick” mark in a CCS directive crashed the CCS table build processor instead of returning an error directive. This was captured in HSTAR 7298. Per an operations request, the SI format was re-configured back to normal at 278/1140Z. The CCS “A” string was re-cycled at 278/1345Z to clear the table build process. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH